# HiMTok: Learning Hierarchical Mask Tokens for Image Segmentation with Large Multimodal Model

### v1

@@ Appendix D Mask Perception
Appendix E Results on general image understanding
@@
We compare methods on MME [19] Perception, VQAv2 [20], and POPE [32]. As shown in Tab. 12, our method is comparable to state-of-the-art LMMs, except in the areas of landmarks, artwork, and OCR, where we do not specifically utilize corresponding training data in this version.
Appendix F More visualizations
@@
Appendix G Prompt design
@@
For the bidirectional information flow between segmentation and grounding, Tabs. 13 and 14 list templates for mask-to-box, and Tabs. 15 and 16 are for coordinate-to-mask. Tab. 17 list templates for segmentation-only responses. Templates in Tab. 18 are used to specify the mask token length from LMM. If visual grounding is the only target without mask tokens, we can refer to Tab. 19.

### v2

@@ Appendix E Mask Perception
Appendix F Results on fine-grained regions
Given that HiMTok decodes masks directly via mask tokens, a natural concern is whether it can maintain high segmentation quality for fine-grained regions without leveraging the fine-grained features of the original image.
Here we show the results on small object segmentation in RefCOCO/+/g. Objects whose mask areas occupy less than 4% of the image are considered as small, resulting in 12.8% samples. Shown in Tab. 12, the cIoUs of ours still have significant priority compared to PSALM [78]. However, the cIoU scores fall significantly behind the overall performance (Tab. 2), which highlights a common challenge.
| Ref COCO | Ref COCO+ | Ref COCOg | ||||||
| val | testA | testB | val | testA | testB | val | test | |
| PSALM [78] | 64.50 | 68.43 | 57.31 | 45.58 | 55.89 | 38.14 | 51.10 | 48.78 |
| ours | 67.15 | 75.00 | 60.34 | 56.81 | 63.06 | 48.29 | 57.71 | 55.99 |
We also review the segmentation boundaries. Bfscore (a boundary-aware F1 metric) on RefCOCO (val) is reported: our method gets 0.927, which is competitive to PSALM (0.936). PSALM integrates Mask2Former that is favored by multi-scale features. We believe there are rooms for future exploration in our paradigm.
Appendix G Results on general image understanding
@@
We compare methods on MME [19] Perception, VQAv2 [20], and POPE [32]. As shown in Tab. 13, our method is comparable to state-of-the-art LMMs, except in the areas of landmarks, artwork, and OCR, where we do not specially utilize corresponding training data in this work. We can conclude that LMMis a comprehensional and general LMM.
Appendix H More visualizations
@@
Appendix I Prompt design
@@
For the bidirectional information flow between segmentation and grounding, Tabs. 14 and 15 list templates for mask-to-box, and Tabs. 16 and 17 are for coordinate-to-mask. Tab. 18 list templates for segmentation-only responses. Templates in Tab. 19 are used to specify the mask token length from LMM. If visual grounding is the only target without mask tokens, we can refer to Tab. 20.